# ERPoT: Effective and Reliable Pose Tracking for Mobile Robots Based on Lightweight and Compact Polygon Maps

### v1

@@ IV-D Quantitative comparative results using public datasets
ERPoT(ours) $\mathbf{0.32}^{\star}$ $\mathbf{136.00}^{\star}$ $\mathbf{18.08}^{\star}$ $\mathbf{21.24}^{\star}$ 19.0319.03{19.03}19.03 $\mathbf{0.88}^{\star}$ $\mathbf{1.23}^{\star}$ $\mathbf{66.18}^{\star}$ $\mathbf{29.99}^{\star}$ $\mathbf{8.68}^{\star}$
@@
ERPoT(ours) $\mathbf{0.32}^{\star}$ $\mathbf{210.54}^{\star}$ $\mathbf{24.92}^{\star}$ $\mathbf{29.47}^{\star}$ $\mathbf{8.05}^{\star}$ $\mathbf{1.03}^{\star}$ $\mathbf{0.34}^{\star}$ $\mathbf{59.30}^{\star}$ $\mathbf{27.73}^{\star}$ $\mathbf{7.07}^{\star}$
@@
On the other hand, the comparative experiments for the Newer College dataset are also carried out, and the polygon map is depicted in Fig. 17. As shown in the detailed map description in the right subfigure, the buildings of the campus environment are represented by multiple compact polygons to obtain the lightweight prior map. Afterward, two test sequences (Newer College-1 and Newer College-2) are used for experimental evaluation of pose tracking, and the quantitative experimental results are shown in Tab. III.
@@
Due to the fact that the Newer College dataset is collected through a handheld platform, the LiDAR data is subject to noise inherent in the movement process. Coupled with the environmental presence of extensive vegetation, these factors collectively pose a significant challenge for the task of pose tracking under pure LiDAR mode. Therefore, for FLOAM_MCL, HDL_localization and SM_MCL, no effective pose tracking outcomes have been achieved in either of the two test sequences. However, our proposed ERPoT maintains successful pose tracking capabilities, even though it exhibits slightly higher translational and rotational errors compared to the previous outcomes. Moreover, the proposed ERPoT also demonstrates the superior performance in terms of runtime.
IV-E Quantitative comparative results using self datasets
Figure 18: Polygon map for Self dataset-01.
Furthermore, to verify the effectiveness of the proposed approach in other complex operating environments, two self-recorded datasets (Self dataset-01 and Self dataset-02) are used for comparative evaluation of pose tracking.
For the Self dataset-01, the corresponding prior polygon map is shown in Fig. 18. There are three test sequences for evaluation of pose tracking, and the comparative experimental results are shown in Tab. IV. Note that the first two test sequences are collected at the same time as the dataset used for constructing the prior map, while the third test sequence (Self dataset-01-3) is obtained a year and a half later. As the detailed comparative experimental results in Tab. IV, compared to the other five approaches, the proposed ERPoT achieves the superior performance in terms of map size, translational error, and rotational error, while simultaneously maintaining high and stable algorithm running efficiency. In particular, the experimental results using Self dataset-01-3 demonstrate that the novel polygon map and pose tracking approach have long-term viability. In contrast, HDL_localization and SM_MCL have encountered failures within this evaluation. The primary reason for these shortcomings is the significant environmental changes that hinder point cloud registration, thereby affecting the stability and accuracy of pose tracking.
@@ IV-E Quantitative comparative results using self datasets
TABLE IV: Experimental results of different approaches using the Self dataset-01
@@
ERPoT(ours) $\mathbf{0.98}^{\star}$ $\mathbf{22.26}^{\star}$ $\mathbf{9.48}^{\star}$ $\mathbf{10.24}^{\star}$ $1.15$ $\mathbf{0.11}^{\star}$ $\mathbf{0.18}^{\star}$ $\mathbf{81.16}^{\star}$ $50.19$ $\mathbf{5.99}^{\star}$
@@
ERPoT(ours) $\mathbf{0.98}^{\star}$ $28.83$ $\mathbf{6.97}^{\star}$ $\mathbf{7.99}^{\star}$ $\mathbf{0.65}^{\star}$ $\mathbf{0.17}^{\star}$ $\mathbf{0.23}^{\star}$ $77.04$ $51.93$ $\mathbf{5.03}^{\star}$
@@
ERPoT(ours) $\mathbf{0.98}^{\star}$ $139.96$ $\mathbf{8.92}^{\star}$ $\mathbf{10.86}^{\star}$ $\mathbf{2.59}^{\star}$ $\mathbf{0.14}^{\star}$ $\mathbf{0.23}^{\star}$ $\mathbf{108.46}^{\star}$ $62.23$ $\mathbf{7.53}^{\star}$
TABLE V: Experimental results of different approaches using the Self dataset-02
[…]
@@
ERPoT(ours) $\mathbf{0.65}^{\star}$ $\mathbf{29.62}^{\star}$ $6.33$ $7.55$ $1.40$ $\mathbf{0.23}^{\star}$ $\mathbf{0.33}^{\star}$ $61.26$ $39.07$ $7.92$
Refer to captionRefer to caption
Figure 19: Polygon map for Self dataset-02.
@@
On the other hand, taking into account the diversity of real-world operating environments, the Self dataset-02 is obtained by selecting an open square that includes uphill and downhill sections. This choice ensures a comprehensive evaluation of the performance of ERPoT across diverse terrains, and simulates the challenges encountered in actual deployments. The obtained prior polygon map is shown in Fig. 19.
@@
More importantly, the comparative experimental results using Self dataset-02 are shown in Tab. V. In particular, the prior map is remarkably compact, with a size of merely $0.65$ MB, and is significantly more concise than the point cloud map and distance field representation. In terms of translational and rotational errors, the proposed ERPoT has achieved optimal or near-optimal performance, all the while maintaining high and stable operational efficiency. Experimental results demonstrate the effectiveness of ERPoT across diverse terrains.
V Conclusion

### v3

@@ IV-D Quantitative comparative results using public datasets
Range_MCL 23.2823.28{23.28}23.28 124.63124.63{124.63}124.63 30.8330.83{30.83}30.83 36.6336.63{36.63}36.63 14.7514.75{14.75}14.75 1.601.60{1.60}1.60 2.052.05{2.05}2.05 $585.20$ $228.98$ $170.01$
ERPoT(ours) $\mathbf{0.32}^{\star}(2749/40341)$ 136.00136.00{136.00}136.00 $\mathbf{18.08}^{\star}$ $\mathbf{21.24}^{\star}$ 19.0319.03{19.03}19.03 $\mathbf{0.88}^{\star}$ $\mathbf{1.23}^{\star}$ $\mathbf{66.18}^{\star}$ $\mathbf{29.99}^{\star}$ $\mathbf{8.68}^{\star}$
@@
Range_MCL 23.2823.28{23.28}23.28 254.92254.92{254.92}254.92 45.7945.79{45.79}45.79 62.4462.44{62.44}62.44 12.8012.80{12.80}12.80 2.012.01{2.01}2.01 2.602.60{2.60}2.60 $585.59$ $422.85$ $158.01$
ERPoT(ours) $\mathbf{0.32}^{\star}(2749/40341)$ $\mathbf{210.54}^{\star}$ $\mathbf{24.92}^{\star}$ $\mathbf{29.47}^{\star}$ $\mathbf{8.05}^{\star}$ $\mathbf{1.03}^{\star}$ $\mathbf{0.34}^{\star}$ $\mathbf{59.30}^{\star}$ $\mathbf{27.73}^{\star}$ $\mathbf{7.07}^{\star}$
@@
On the other hand, comparative experiments on the Newer College dataset are conducted, with the polygon map shown in Fig. 18. The campus buildings are represented by compact polygons to create a lightweight prior map, as detailed in the right subfigure. Two test sequences (Newer College-1 and Newer College-2) are evaluated for pose tracking, with quantitative results presented in Tab. V.
@@
The Newer College dataset, collected via a handheld platform, contains noisy LiDAR data due to movement. Combined with extensive vegetation, these factors create significant challenges for pure LiDAR pose tracking. As a result, methods like FLOAM_MCL, HDL_localization, and SM_MCL fail to achieve effective tracking in both test sequences. However, our proposed ERPoT maintains successful pose tracking capabilities, even though it exhibits slightly higher translational and rotational errors. Moreover, the proposed ERPoT also demonstrates the superior performance in terms of runtime.
(a) Portable-campus
(b) Portable-garden
(c) Portable-canteen
(d) Portable-building
Figure 19: Polygon maps for the FusionPortable dataset.
To further validate the effectiveness and reliability of ERPoT, additional comparative experiments are conducted using the Portable-campus and Portable-building datasets. The polygon maps for these scenarios are shown in Fig. 19(a) and Fig. 19(d), with results detailed in Tab. VI. Notably, the prior map size ranges from $0.1$ to $0.2$ MB, demonstrating remarkable performance. ERPoT also shows superior performance in translational error, rotational error, and runtime.
@@
TABLE VI: Experimental results of different approaches using the FusionPortable dataset(outdoor)
[…]
Portable-campus ALOAM_MCL 24.7724.77{24.77}24.77 $206.92$ $45.65$ $57.48$ $11.51$ $1.62$ $2.15$ $590.09$ $266.80$ $134.53$
FLOAM_MCL $197.48$ $65.63$ $75.57$ $16.05$ $2.18$ $3.14$ $292.33$ $144.75$ $82.73$
KISS_MCL $132.39$ $61.39$ $70.56$ $7.16$ $1.40$ $1.78$ $390.19$ $116.33$ $49.14$
HDL_localization $249.66$ $25.84$ $31.02$ $12.13$ $1.12$ $1.77$ $564.45$ $127.97$ $126.70$
SM_MCL $2369.31$ $256.00$ $35.08$ $44.02$ $\mathbf{5.85}^{\star}$ $\mathbf{0.96}^{\star}$ $1.48$ $\mathbf{71.43}^{\star}$ $\mathbf{32.71}^{\star}$ $\mathbf{5.48}^{\star}$
Range_MCL 10.6810.68{10.68}10.68 80.6680.66{80.66}80.66 34.8234.82{34.82}34.82 38.3038.30{38.30}38.30 19.5919.59{19.59}19.59 5.575.57{5.57}5.57 3.593.59{3.59}3.59 $441.12$ $239.27$ $151.00$
ERPoT(ours) $\mathbf{0.10}^{\star}(960/12655)$ $\mathbf{42.38}^{\star}$ $\mathbf{17.08}^{\star}$ $\mathbf{19.30}^{\star}$ 6.606.60{6.60}6.60 1.111.11{1.11}1.11 $\mathbf{1.47}^{\star}$ 101.89101.89{101.89}101.89 58.7658.76{58.76}58.76 10.5110.51{10.51}10.51
Portable-building-1 ALOAM_MCL 42.3342.33{42.33}42.33 $211.11$ $55.91$ $80.42$ $17.46$ $2.54$ $3.56$ $564.36$ $257.12$ $129.66$
FLOAM_MCL $82.48$ $26.82$ $32.35$ $6.91$ $1.56$ $2.00$ $394.81$ $123.75$ $25.10$
KISS_MCL $100.37$ $31.07$ $39.66$ 6.886.88{6.88}6.88 1.741.74{1.74}1.74 $2.23$ $185.86$ $63.45$ $29.50$
HDL_localization -- -- -- -- -- -- -- -- --
SM_MCL $3335.99$ $147.95$ $34.63$ $42.14$ $7.37$ $1.22$ $1.70$ $\mathbf{52.08}^{\star}$ $\mathbf{28.30}^{\star}$ $7.29$
Range_MCL 10.1510.15{10.15}10.15 103.09103.09{103.09}103.09 41.8541.85{41.85}41.85 50.3250.32{50.32}50.32 12.7012.70{12.70}12.70 1.851.85{1.85}1.85 2.692.69{2.69}2.69 $100.65$ $47.44$ $8.61$
ERPoT(ours) $\mathbf{0.23}^{\star}(2290/30488)$ $\mathbf{49.78}^{\star}$ $\mathbf{12.53}^{\star}$ $\mathbf{15.10}^{\star}$ $\mathbf{5.33}^{\star}$ $\mathbf{0.46}^{\star}$ $\mathbf{0.63}^{\star}$ 81.0381.03{81.03}81.03 56.8156.81{56.81}56.81 $\mathbf{5.33}^{\star}$
Portable-building-2 ALOAM_MCL ${42.33}$- -- -- -- -- -- -- -- --
FLOAM_MCL $78.11$ $25.02$ $29.36$ $10.58$ $2.03$ $2.63$ $216.99$ $115.48$ $44.75$
KISS_MCL $78.40$ $24.06$ $28.42$ 14.4814.48{14.48}14.48 2.142.14{2.14}2.14 $3.01$ $194.57$ $69.01$ $32.71$
HDL_localization -- -- -- -- -- -- -- -- --
SM_MCL $3335.99$- -- -- -- -- -- -- -- --
Range_MCL 10.1510.15{10.15}10.15 187.95187.95{187.95}187.95 53.3153.31{53.31}53.31 73.6173.61{73.61}73.61 13.0713.07{13.07}13.07 2.512.51{2.51}2.51 3.403.40{3.40}3.40 $347.70$ $259.41$ $130.56$
ERPoT(ours) $\mathbf{0.23}^{\star}(2290/30488)$ $\mathbf{38.15}^{\star}$ $\mathbf{11.92}^{\star}$ $\mathbf{13.06}^{\star}$ $\mathbf{3.06}^{\star}$ $\mathbf{0.61}^{\star}$ $\mathbf{0.79}^{\star}$ $\mathbf{95.26}^{\star}$ $\mathbf{59.26}^{\star}$ $\mathbf{9.23}^{\star}$
IV-E Comparative experiments for indoor environments
TABLE VII: Experimental results of different approaches using the FusionPortable dataset(indoor)
[…]
Portable-garden ALOAM_MCL ${10.53}$- -- -- -- -- -- -- -- --
FLOAM_MCL $99.01$ $20.97$ $23.64$ $10.76$ $1.52$ $2.02$ $203.81$ $122.94$ $10.76$
KISS_MCL $168.52$ $18.85$ $26.65$ 16.0316.03{16.03}16.03 1.261.26{1.26}1.26 $1.74$ $159.38$ $65.91$ $17.54$
HDL_localization -- -- -- -- -- -- -- -- --
SM_MCL $644.31$- -- -- -- -- -- -- -- --
Range_MCL 16.1116.11{16.11}16.11 $\mathbf{46.74}^{\star}$ 20.1320.13{20.13}20.13 21.8521.85{21.85}21.85 8.328.32{8.32}8.32 1.381.38{1.38}1.38 1.801.80{1.80}1.80 $485.29$ $398.88$ $166.04$
ERPoT(ours) $\mathbf{0.06}^{\star}(627/8135)$ 73.0373.03{73.03}73.03 $\mathbf{15.39}^{\star}$ $\mathbf{17.72}^{\star}$ $\mathbf{4.31}^{\star}$ $\mathbf{0.66}^{\star}$ $\mathbf{0.90}^{\star}$ $\mathbf{103.98}^{\star}$ $\mathbf{64.34}^{\star}$ $\mathbf{9.45}^{\star}$
Portable-canteen ALOAM_MCL ${13.72}$- -- -- -- -- -- -- -- --
FLOAM_MCL $81.12$ $19.09$ $23.10$ $9.05$ $1.49$ $1.95$ $322.74$ $113.17$ $17.34$
KISS_MCL $71.18$ $18.98$ $24.29$ $5.78$ $1.46$ $1.82$ $154.35$ $\mathbf{55.63}^{\star}$ $13.18$
HDL_localization -- -- -- -- -- -- -- -- --
SM_MCL $885.53$- -- -- -- -- -- -- -- --
Range_MCL 18.4618.46{18.46}18.46 $\mathbf{37.39}^{\star}$ 13.3213.32{13.32}13.32 14.9314.93{14.93}14.93 8.198.19{8.19}8.19 1.311.31{1.31}1.31 1.731.73{1.73}1.73 $588.19$ $486.95$ $82.51$
ERPoT(ours) $\mathbf{0.11}^{\star}(1115/15195)$ 43.0543.05{43.05}43.05 $\mathbf{9.17}^{\star}$ $\mathbf{10.22}^{\star}$ $\mathbf{5.02}^{\star}$ $\mathbf{0.56}^{\star}$ $\mathbf{0.79}^{\star}$ $\mathbf{99.08}^{\star}$ 62.4562.45{62.45}62.45 $\mathbf{11.16}^{\star}$
To further validate the effectiveness of ERPoT in indoor environments, comparative experiments are conducted using the Portable-garden and Portable-canteen datasets. These two scenarios represent typical indoor garden and canteen environments, respectively, and include dynamic pedestrian information. Detailed information is presented in Fig. 11(c).
The prior polygon maps are generated from the nighttime LiDAR sequences, as shown in Fig. 19(b) and Fig. 19(c). Notably, the prior map size of ERPoT remains exceptionally compact, even under $0.1$MB in the garden environment. Comparative results are presented in Tab. VII, highlighting the challenges of pose tracking in complex indoor settings, which lead to the failure of some methods like ALOAM_MCL, HDL_localization, and SM_MCL. Despite these challenges, ERPoT demonstrates excellent performance, particularly in runtime, achieving stable results with a 128-line LiDAR.
IV-F Quantitative comparative results using self datasets
(a) Self dataset-01
(b) Self dataset-02
Figure 20: Polygon maps for the Self dataset.
Furthermore, to verify the effectiveness of the proposed approach in other complex operating environments, two self-recorded datasets (Self dataset-01 and Self dataset-02) are utilized for comparative evaluation of pose tracking. For Self dataset-01 and Self dataset-02, the ground truth is derived using LIO-SAM [5], which integrates data from the LiDAR sensor, high-precision GPS, and an inertial measurement unit (IMU). This integration yields the ground truth for keyframes and records the relative pose of each frame with respect to its corresponding keyframe. Finally, linear interpolation is applied to calculate the pose of each frame, thereby obtaining the ground truth for the entire dataset.
For the Self dataset-01, the corresponding prior polygon map is shown in Fig. 20(a). There are three test sequences for evaluation of pose tracking, and the comparative experimental results are shown in Tab. VIII. Note that the first two test sequences are collected at the same time as the dataset used for constructing the prior map, while the third test sequence (Self dataset-01-3) is obtained a year and a half later. As the detailed comparative experimental results in Tab. VIII, compared to the other six approaches, the proposed ERPoT achieves the superior performance in terms of map size, translational error, and rotational error, while simultaneously maintaining high and stable algorithm running efficiency. In particular, the experimental results using Self dataset-01-3 demonstrate that the novel polygon map and pose tracking approach have long-term viability, and the long-term environmental changes in the testing environment are illustrated in Fig. 21. In contrast, HDL_localization and SM_MCL have encountered failures within this evaluation. Unlike ALOAM_MCL, FLOAM_MCL, and KISS_MCL, these methods rely solely on the prior point cloud map obtained over a year and a half ago. The primary reason for these shortcomings is the significant environmental changes that hinder point cloud registration, thereby affecting the stability and accuracy of pose tracking.
Figure 21: Diagram of point cloud changes in the testing environment for Self dataset-01. The colored point cloud represents the prior map obtained a year and a half ago, while the black point cloud represents the point cloud information of the current testing environment.
TABLE VIII: Experimental results of different approaches using the Self dataset
@@ IV-F Quantitative comparative results using self datasets
Range_MCL 56.1556.15{56.15}56.15 73.2673.26{73.26}73.26 29.8529.85{29.85}29.85 32.4432.44{32.44}32.44 9.009.00{9.00}9.00 0.880.88{0.88}0.88 1.621.62{1.62}1.62 $600.22$ $537.70$ $127.39$
ERPoT(ours) $\mathbf{0.98}^{\star}(8619/122066)$ $\mathbf{22.26}^{\star}$ $\mathbf{9.48}^{\star}$ $\mathbf{10.24}^{\star}$ $1.15$ $\mathbf{0.11}^{\star}$ $\mathbf{0.18}^{\star}$ $\mathbf{81.16}^{\star}$ $50.19$ $\mathbf{5.99}^{\star}$
@@
Range_MCL 56.1556.15{56.15}56.15 73.2373.23{73.23}73.23 29.9629.96{29.96}29.96 30.8730.87{30.87}30.87 2.732.73{2.73}2.73 0.700.70{0.70}0.70 0.890.89{0.89}0.89 $610.77$ $468.34$ $63.05$
ERPoT(ours) $\mathbf{0.98}^{\star}(8619/122066)$ $28.83$ $\mathbf{6.97}^{\star}$ $\mathbf{7.99}^{\star}$ $\mathbf{0.65}^{\star}$ $\mathbf{0.17}^{\star}$ $\mathbf{0.23}^{\star}$ $77.04$ $51.93$ $\mathbf{5.03}^{\star}$
@@
Range_MCL 56.1556.15{56.15}56.15 301.81301.81{301.81}301.81 37.3437.34{37.34}37.34 52.8352.83{52.83}52.83 26.7626.76{26.76}26.76 1.561.56{1.56}1.56 3.393.39{3.39}3.39 $724.46$ $421.64$ $210.58$
ERPoT(ours) $\mathbf{0.98}^{\star}(8619/122066)$ $139.96$ $\mathbf{8.92}^{\star}$ $\mathbf{10.86}^{\star}$ $\mathbf{2.59}^{\star}$ $\mathbf{0.14}^{\star}$ $\mathbf{0.23}^{\star}$ $\mathbf{108.46}^{\star}$ $\mathbf{62.23}^{\star}$ $\mathbf{7.53}^{\star}$
@@
Range_MCL 28.8328.83{28.83}28.83 350.34350.34{350.34}350.34 131.15131.15{131.15}131.15 154.27154.27{154.27}154.27 10.3910.39{10.39}10.39 1.561.56{1.56}1.56 2.412.41{2.41}2.41 $803.37$ $457.29$ $209.51$
ERPoT(ours) $\mathbf{0.65}^{\star}(5855/78864)$ $\mathbf{29.62}^{\star}$ $6.33$ $7.55$ $1.40$ $\mathbf{0.23}^{\star}$ $\mathbf{0.33}^{\star}$ $61.26$ $39.07$ $7.92$
@@
On the other hand, taking into account the diversity of real-world operating environments, the Self dataset-02 is obtained by selecting an open square that includes uphill and downhill sections. This choice ensures a comprehensive evaluation of the performance of ERPoT across diverse terrains, and simulates the challenges encountered in actual deployments. The obtained prior polygon map is shown in Fig. 20(b).
@@
More importantly, the comparative experimental results using Self dataset-02 are shown in Tab. VIII. In particular, the prior map is remarkably compact, with a size of merely $0.65$ MB, and is significantly more concise than the point cloud map and distance field representation. In terms of translational and rotational errors, the proposed ERPoT has achieved optimal or near-optimal performance, while maintaining high and stable operational efficiency. Experimental results demonstrate the effectiveness of ERPoT across diverse terrains.
IV-G Ablation study
TABLE IX: Experimental results of ablation study for ERPoT without(w/o) point-vertex constraint or point-edge constraint
Dataset No. w/o Point-vertex constrait w/o Point-edge constrait
Succ. Impact on pose tracking Succ. Impact on pose tracking
KITTI 00-1 63.2263.22{63.22}63.22 $11.52$ $14.79$ 48.8848.88{48.88}48.88 16.1016.10{16.10}16.10 19.0619.06{19.06}19.06
0.840.84{0.84}0.84 $0.37$ $0.42$ 0.870.87{0.87}0.87 0.400.40{0.40}0.40 0.450.45{0.45}0.45
00-2 $20.87$ $7.68$ $8.77$ 33.1733.17{33.17}33.17 10.7810.78{10.78}10.78 12.2112.21{12.21}12.21
$0.83$ $0.21$ $0.26$ 1.281.28{1.28}1.28 0.240.24{0.24}0.24 0.300.30{0.30}0.30
00-3 43.3843.38{43.38}43.38 13.8913.89{13.89}13.89 18.2518.25{18.25}18.25 38.1938.19{38.19}38.19 15.9515.95{15.95}15.95 19.3719.37{19.37}19.37
0.580.58{0.58}0.58 0.160.16{0.16}0.16 18.2518.25{18.25}18.25 0.550.55{0.55}0.55 0.160.16{0.16}0.16 0.200.20{0.20}0.20
02-1 35.5435.54{35.54}35.54 12.3012.30{12.30}12.30 14.1114.11{14.11}14.11 30.3630.36{30.36}30.36 12.0312.03{12.03}12.03 13.7513.75{13.75}13.75
0.500.50{0.50}0.50 0.250.25{0.25}0.25 0.280.28{0.28}0.28 0.700.70{0.70}0.70 0.230.23{0.23}0.23 0.260.26{0.26}0.26
02-2 86.7986.79{86.79}86.79 19.3119.31{19.31}19.31 24.2224.22{24.22}24.22 77.2577.25{77.25}77.25 19.1519.15{19.15}19.15 22.5422.54{22.54}22.54
0.310.31{0.31}0.31 0.390.39{0.39}0.39 24.2224.22{24.22}24.22 0.860.86{0.86}0.86 0.290.29{0.29}0.29 0.350.35{0.35}0.35
05-1 29.8329.83{29.83}29.83 7.307.30{7.30}7.30 8.678.67{8.67}8.67 28.3828.38{28.38}28.38 11.0711.07{11.07}11.07 12.2012.20{12.20}12.20
1.461.46{1.46}1.46 8.678.67{8.67}8.67 0.310.31{0.31}0.31 1.581.58{1.58}1.58 0.260.26{0.26}0.26 0.360.36{0.36}0.36
05-2 42.8042.80{42.80}42.80 8.548.54{8.54}8.54 10.1110.11{10.11}10.11 36.8036.80{36.80}36.80 14.3314.33{14.33}14.33 15.7615.76{15.76}15.76
0.760.76{0.76}0.76 0.250.25{0.25}0.25 0.290.29{0.29}0.29 0.770.77{0.77}0.77 0.280.28{0.28}0.28 0.320.32{0.32}0.32
06-1 44.9544.95{44.95}44.95 7.207.20{7.20}7.20 8.398.39{8.39}8.39 26.3126.31{26.31}26.31 9.069.06{9.06}9.06 9.979.97{9.97}9.97
0.570.57{0.57}0.57 0.240.24{0.24}0.24 0.260.26{0.26}0.26 0.520.52{0.52}0.52 0.230.23{0.23}0.23 0.260.26{0.26}0.26
08-1 72.3672.36{72.36}72.36 17.8817.88{17.88}17.88 23.5723.57{23.57}23.57
1.291.29{1.29}1.29 0.330.33{0.33}0.33 0.400.40{0.40}0.40
Newer College 00-1 146.41146.41{146.41}146.41 19.5619.56{19.56}19.56 22.7022.70{22.70}22.70
$18.20$ $0.87$ $1.19$
00-2
FusionPortable campus 46.2346.23{46.23}46.23 46.2346.23{46.23}46.23 20.4720.47{20.47}20.47
$6.58$ $1.01$ $1.39$
building-1 51.6651.66{51.66}51.66 14.4114.41{14.41}14.41 16.9016.90{16.90}16.90
5.595.59{5.59}5.59 0.460.46{0.46}0.46 0.640.64{0.64}0.64
building-2 46.2746.27{46.27}46.27 13.6913.69{13.69}13.69 14.9314.93{14.93}14.93
3.153.15{3.15}3.15 0.630.63{0.63}0.63 0.810.81{0.81}0.81
garden 266.39266.39{266.39}266.39 17.0317.03{17.03}17.03 27.7227.72{27.72}27.72
21.9021.90{21.90}21.90 0.740.74{0.74}0.74 1.481.48{1.48}1.48
canteen 47.0047.00{47.00}47.00 11.3211.32{11.32}11.32 12.5312.53{12.53}12.53
5.265.26{5.26}5.26 0.610.61{0.61}0.61 0.850.85{0.85}0.85
Self dataset 01-1 $20.29$ $7.36$ $7.86$ 24.7024.70{24.70}24.70 10.9510.95{10.95}10.95 11.8911.89{11.89}11.89
1.511.51{1.51}1.51 $0.10$ $0.18$ $1.13$ 0.120.12{0.12}0.12 0.190.19{0.19}0.19
01-2 159.53159.53{159.53}159.53 6.176.17{6.17}6.17 14.2414.24{14.24}14.24 71.7371.73{71.73}71.73 9.429.42{9.42}9.42 11.0811.08{11.08}11.08
0.850.85{0.85}0.85 0.170.17{0.17}0.17 0.220.22{0.22}0.22 0.680.68{0.68}0.68 0.180.18{0.18}0.18 0.240.24{0.24}0.24
01-3 106.83106.83{106.83}106.83 10.1610.16{10.16}10.16 12.2012.20{12.20}12.20
2.552.55{2.55}2.55 0.150.15{0.15}0.15 0.240.24{0.24}0.24
02-1 31.3931.39{31.39}31.39 $7.44$ 8.598.59{8.59}8.59
$1.13$ $0.23$ $0.33$
To systematically evaluate the contributions of various components in our proposed ERPoT, an ablation study is conducted. This study aims to provide insights into how each component affects the overall performance of the system, particularly focusing on the point-vertex and point-edge constraints. Importantly, we evaluate the performance of ERPoT both with and without these constraints across all datasets.
The results of the ablation study are presented in Tab. IX. The top three columns represent translational errors, while the bottom three columns show rotational errors. Results that are worse than those of the complete ERPoT are highlighted with gray shading. It is clear that the point-edge constraint plays a significant role in successful pose tracking, and incorporating the point-edge constraint can further enhance the accuracy.
V Discussion
V-A Limitations and future work
While ERPoT achieves excellent performance in map size, translational error, rotational error, and runtime, it has limitations. Specifically, ERPoT relies on ground segmentation to transform dense 3D LiDAR point clouds into sparse 2D scans. This dependency can cause issues in environments where reliable ground information cannot be extracted, such as in scenarios with uneven terrain or cluttered scenes with low obstacles. In these cases, the ground segmentation process may fail, leading to degraded performance.
Moreover, the dependency of ERPoT on ground segmentation makes it less versatile compared to other pose-tracking methods that do not require such preprocessing steps. To address these limitations, future work could explore alternative preprocessing techniques that are more robust to environmental variations. On the other hand, it is worth noting that there are alternative approaches (e.g., ROS packages like pointcloud_to_laserscan) that can extract a 2D scan from a 3D LiDAR without relying on ground detection. Additionally, the lightweight and compact polygon map is ideal for multi-robot collaborative tasks, enabling efficient map sharing, faster communication, and enhanced system performance, while its concise representation supports rapid scene recognition, crucial for real-time applications like search and rescue or dynamic industrial settings.
V-B Probabilistic model extension
In the field of robot pose tracking, Monte Carlo localization-based approaches play the crucial role, which makes use of a probabilistic estimation technique to properly handle uncertainties about initial, motion process, and sensor observation. In this section, we explore the integration of our proposed approach as an measurement model within the Monte Carlo Localization (MCL) framework. Our proposed ERPoT is a approach which relies on a lightweight and compact polygon map for pose tracking. Specifically, it converts dense 3D LiDAR point clouds into sparse 2D scans and employs a novel cost function for pose estimation through point-polygon matching. This approach ensures efficient and accurate pose tracking, making it suitable for integration with MCL.
The complete probabilistic localization process contains initialization, motion model, measurement model, resampling and pose estimation. In this section, we focus on the measurement model, which makes use of the kernel contributions on the proposed novel polygon map representation and the data association between the point and the polygon. For time stamp $t$, given the generated sparse 2D scan observation $\mathbf{C}_{t}$ and the lightweight polygon map ${\cal M}_{p}$, the measurement model is represented as follows
$p(\mathbf{C}_{t}^{hit}|\tilde{\mathbf{s}}_{t},{\cal M}_{p})=\prod\nolimits_{i=1}^{K}{p\left({\mathbf{p}_{i}^{2D}}|\tilde{\mathbf{s}}_{t},\mathbf{P}_{i\to j}\right)},$ (35)
where the $j$-th corresponding polygon $\mathbf{P}_{j}\in{\cal M}_{p}$, and $\mathbf{C}_{t}^{hit}$ is subset of $\mathbf{C}_{t}$ with $K$ scan points that meet the condition
$d^{*}(\mathbf{p}_{i}^{2D},{\cal M}_{p})={\min_{j=1,2,...,m}}d\left({\mathbf{p}_{i}^{2D},\mathbf{P}_{j}}\right)<\delta_{d},$ (36)
and $\delta_{d}$ is the data association threshold. Accordingly,
$p\left({\mathbf{p}_{i}^{2D}}|\tilde{\mathbf{s}}_{t},\mathbf{P}_{i\to j}\right)=\mathcal{N}\left(d^{*};0,\sigma_{hit}^{2}\right),$ (37)
where $\sigma_{hit}$ is the standard deviation.
VI Conclusion